A sum of money at simple interest amounted to P7800 in 5 years and to P8840 in 9 years. Find the principal and the rate.
Answer by mananth(16946)

let invested amount be x
interest = 7800-x for 5 years
for 1 year = 7800-x /5
...
for 9 years interest = 8840-x
for 1 year = 8840-x/9
.....
8840-x/9 = 7800-x / 5
5(8840-x)=9(7800-x)
44200-5x=69600-9x
9x-5x=69600-44200
4x=15400
x=3850 investment
Interest for 1 year = 7800-3850 /5
=3950/5
=790
3850 790
100 ?
79000/3850
20.5% =rate